OpenStudy (anonymous):

PLEASE HELP, I GOT A FINAL EXAM... in 3 weeks :) A math teacher is randomly distributing 15 black dry erase markers and 6 blue dry erase markers. What is the probability that the first marker handed out will be blue and the second marker will be black?

OpenStudy (kropot72):

There are 21 markers in total. The probability of handing out a blue marker first is 6/21. If a blue marker is the first to be handed out, there are 20 markers remaining of which15 are black. Therefore the probability of a black marker being the second to be handed out is 15/20. the required probability is therefore given by: \[\large \frac{6}{21} \times \frac{15}{20}=you\ can\ calculate\]

OpenStudy (kropot72):

@Romwil Do you understand?

OpenStudy (anonymous):

I'm stupid I got that, but I did 15/21 * 6/21 I didn't read lol thx I owe you one

OpenStudy (kropot72):

You're welcome :)